# Finite Sample Analysis of Distribution-Free Confidence Ellipsoids for Linear Regression

### v1

@@ I Introduction
Building mathematical models of unknown systems based on noisy empirical data is a fundamental problem in system identification, signal processing and machine learning. There are several standard methods that provide point estimates with uncertainty quantification based on limiting distributions. Typical examples include prediction error methods [1] with asymptotic confidence regions derived from the central limit theorem. A considerable disadvantage of these confidence sets is that they are only approximately correct for finite samples. Consequently, their application could be problematic in domains with strong stability or safety requirements.
@@
In order to overcome this issue, a substantial amount of the recent research in system identification focus on approaches with non-asymptotic guarantees [2]. One of the most common techniques to derive probability approximately correct (PAC) bounds for an estimate, such as LS, is based on the theory of concentration inequalities, which induces confidence regions.
@@
Several works analyzed (closed-loop) dynamical systems, where the sample complexity of transition matrix estimations were studied for fully observable [3] and also non-observable state space settings [4]. A signal processing focused approach was investigated in [5], where a low-rank and structured sparse high-dimensional vector autoregressive (VAR) system was analyzed, and a new estimation technique was proposed with non-asymptotic probabilistic upper bounds on the estimation error. In the aforementioned works, there are strong assumptions on the noises, namely that they follow some specific distribution, in most cases they are Gaussian. Another limitation is that the confidence regions heavily rely on specific constants (hyper-parameters), which are coming from the assumptions on the system setting and the distributions of the inputs and the outputs. In practice, these hyper-parameters are usually unknown and only estimates can be used or bounds based on domain specific knowledge of experts. For finite impulse response (FIR) systems, PAC bounds for the LS estimate were given in [6], under a centered subgaussian noise assumption. Despite the relaxed noise assumption compared to the Gaussian case, building confidence regions using this results would still rely on the variance proxy of the subgaussian distribution.
@@
Theoretical guarantees besides exact coverage were rigorously proven for SPS, such as uniform strong consistency[19], which means that the SPS regions almost surely shrink around the true parameters, as the sample size increases. This result however is asymptotic, and the finite sample performance of SPS remained an open question until recently. In our previous work [20], we have investigated the sample complexity of SPS for linear regression problems, i.e., we have derived non-asymptotic PAC bounds for the volumes of SPS regions. We have also shown that the sizes of SPS regions decrease at an optimal rate $\mathcal{O}(1/\sqrt{n})$, where $n$ is the sample size.
All of the theoretical results mentioned above considers the fundamental construction of SPS: the indicator function. The standard SPS, given by its indicator function, is a hypothesis test that checks whether a given parameter is included in the confidence region. Using the indicator function to determine all the points of the confidence set would be computationally demanding, therefore in [14] an ellipsoidal outer approximation (EOA) algorithm was proposed, which gives a compact representation of the SPS region. An important feature of the constructed ellipsoid is that its center and shape matrix coincide with that of the confidence ellipsoid constructed using the classical asymptotic theory, only its radius is different. This radius, which ensures finite sample guarantees, can be computed in polynomial time, by solving semidefinite optimization problems. The EOA construction of SPS was later generalized to ARX [21] and closed-loop state space models [18], as well.
In this we work provide a finite sample analysis of the EOA of SPS for general linear regression problems with exogenous regressors, e.g., for FIR systems. We prove high probability upper bounds on the sizes of the resulting confidence ellipsoids. Our analysis builds on some results from the sample complexity analysis of standard SPS regions [20], however, we analyse the finite sample properties of a convex semidefinite program, which requires a significantly different approach than deriving concentration inequalities for the indicator function. We also show that the obtained bounds are “close” to the bounds of the original SPS region and that they decrease at the optimal rate, that is, $\mathcal{O}(1/\sqrt{n})$. Simulation experiments that demonstrate the difference between our theoretical bounds and the empirical sizes of the regions are also presented. We emphasize that although our analysis use distribution and regressor dependent parameters, the SPS EOA does not require these parameters, and it can be applied under very general assumptions, unlike standard non-asymptotic solutions, for example, based on PAC bounds or set membership ellipsoids.
@@ II-B Assumptions
An (integrable) random variable $W$ is $\sigma$-subgaussian[22, Definition 2.2], if for all $\lambda\in\mathbb{R},$ there is a $\sigma>0$, such that
@@
Note that the quantity $\sigma^{2}$ is often referred to as the variance proxy. Naturally, any Gaussian random variable with variance $\sigma^{2}$ is $\sigma$-subgaussian. Furthermore, any distribution with a bounded support on $[a,b]$ is subgaussian with $\sigma=(b-a)/2$ [22, Example 2.4]. Standard examples of such distributions include the uniform, the triangular and the beta distributions.
@@
A4 applies the definition of coherence from [23, Definition 1.2] along with the facts that $\text{range}(\Phi_{n})=\text{range}(\Phi_{{\scriptscriptstyle Q},n})$ and $\Phi_{{\scriptscriptstyle Q},n}\Phi_{{\scriptscriptstyle Q},n}^{\mathrm{T}}$ is an orthogonal projection onto $\text{range}(\Phi_{n})$. A consequence of this is that $1\leq\mu(\Phi_{n})\leq n/d$ [23]. This assumption provides an upper bound on how much of the excitation can be “concentrated” to a single time step, i.e., onto a particular regressor $\varphi_{t}$. In the presented theorems A4 will be assumed alongside with A2, hence, the thin QR-decomposition of $\Phi_{n}=\Phi_{{\scriptscriptstyle Q},n}\Phi_{{\scriptscriptstyle R},n}$ is unique, since matrix $\Phi_{{\scriptscriptstyle R},n}$ is full rank.
@@
The SPS EOA algorithm, presented as Algorithm 3, requires much milder assumptions on the noises and regressors. The introduced stronger assumptions are needed to analyse the size of the constructed confidence ellipsoids for any finite sample size and almost all realizations of the regressor vectors.
@@ III-A The SPS-Indicator algorithm
$f(\delta)\,\doteq\,\begin{cases}\sigma\hskip 0.85358pt(\hskip 0.85358pt8\hskip 1.42262ptd\ln^{\frac{1}{2}}(\tfrac{4}{\delta})+d)^{\frac{1}{2}}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 2,\\[2.84526pt] \sigma\left(8\ln(\tfrac{4}{\delta})+d\right)^{\frac{1}{2}}&0<\delta<4e^{-(nd\lambda_{0})^{2}},\\ \end{cases}$
@@ IV Sample complexity of the
The SPS ellipsoid gives a compact representation of the confidence region, and compared to applying the SPS-Indicator, it can be much more easily used in various applications, from robust optimization to risk management. It is also significantly less demanding from a computational point of view, as it can be computed in polynomial time. The following theorem formalizes the sample complexity of these ellipsoids, i.e., it shows how their radii shrink as the sample size increases.
@@
By expanding $\left\lVert S_{0}(\theta)\right\rVert^{2}$ as in [14, Section VI.], we get that
@@
Them, by introducing $z\doteq\bar{R}_{n}^{\frac{1}{2}}(\theta-\hat{\theta}_{n})$, the optimization problem (16) can be reformulated as [14, Section VI.]
@@
Note that in [25, Theorem 1] it has been shown that the SPS confidence regions are bounded, if both the perturbed and the unperturbed regressors span the whole space. This means that assuming A2 and $m=2$, the SPS region is bounded if $A$ is positive definite, which is independent of $\{W_{t}\}$.
In case of a bounded SPS confidence region, the ellipsoid $\mathcal{E}_{1}=\{\theta:\|S_{0}(\theta)\|^{2}-\|S_{1}(\theta)\|^{2}\leq 0\}$ is compact and non-empty, since it is a closure of the bounded SPS-Indicator set $\{\theta:\|S_{0}(\theta)\|^{2}\prec_{\pi}\|S_{1}(\theta)\|^{2}\}$, and by construction of the SPS region $\|S_{0}(\hat{\theta}_{n})\|^{2}=0$, therefore the LSE is always included in $\mathcal{E}_{1}$. In Lemma 4 we show that if the SPS region is bounded, given $\{\varphi_{t}\}$ and $\{\alpha_{1,t}\}$, it almost surely holds that $\|S_{1}(\hat{\theta}_{n})\|^{2}\neq 0$. From the previous two properties, it follows that the optimization problem (16) is strictly feasible, i.e., Slater’s condition hold, if we work with bounded SPS confidence regions. Consequently, strong duality holds [26, Appendix B], therefore the value of the above optimization program (19) is equal to the value of its dual which can be formulated by using the Schur complement, as follows:
@@
The semidefinite optimization problem (27) is convex and it can be solved using standard convex optimization tools.
@@
$\gamma^{*}\doteq\begin{cases}\gamma_{0}^{*}&\text{if the SPS region is bounded}\\ \infty&\text{if the SPS region is unbounded}\\ \end{cases}$ (28)
@@
Lower bounds on variables $\gamma_{0}$ and $\xi$ (27) can be determined by the conditions for positive semidefiniteness of (generalized) Schur complements [26, Appendix A.5.5], that is
@@
To give a lower bound on $\xi$, we will reformulate the condition $\xi A^{\prime}_{0}(\xi)\succeq 0$. First, we rewrite $A$ the following way
@@
The proof of Lemma 3 is built upon the same ideas as the proof of [20, Lemma 4], but since our second term is different, we present it in Appendix A, for the sake of completeness.
@@
Let us fix an index $i\in[m-1],$ and introduce the “bad” event $\mathcal{B}_{i}$ that the $i$th ellipsoid is larger than our bound, i.e.,
@@
From Lemma 3, we know that $\mathbb{P}(\mathcal{B}_{i})\leq\delta$, for all $i$. Then, by using a similar argument as in the proof of [24, Theorem 3], we can write the overall bad event, i.e., that there are at least $m-q$ ellipsoids for which our bound does not hold, as
@@
V Simulation Experiments

### v2

@@ I Introduction
Building mathematical models of unknown systems based on noisy empirical data is a fundamental problem in system identification, signal processing, machine learning and statistics. Linear stochastic systems form an important class of models [1, 2], not only because they are better-understood theoretically and can be handled more easily, but they are also widely used in practical applications. Moreover, they can serve as a stepping stone to understanding more complex models, such as nonlinear systems, which could be handled by linearization around a working point [3].
@@
Linear regression is the most fundamental formulation of linear estimation problems. There are several standard methods that provide point estimates for this problem with uncertainty quantification based on limiting distributions. Typical examples include the least-squares (LS) estimate with asymptotic confidence regions derived from the central limit theorem [4]. A considerable disadvantage of these confidence sets is that they are only approximately correct for finite samples. Consequently, their application could be problematic in domains where robustness is an important factor or strong stability requirements are posed. Furthermore, even in not safety critical applications, the decision making process based on the estimated model could be drastically improved by taking non-asymptotically guaranteed uncertainties into account.
Robust statistical methods in linear estimation are crucial for signal processing applications [5]. Due to their connections to sparse signal recovery and compressed sensing, robust linear regression methods [6, 7, 8] are actively studied by the signal processing community. These methods offer estimation techniques that are more robust than LS, allowing for more effective outlier management. We would argue that probabilistic uncertainty quantification for the LS estimate in this case could also be a possible direction. Robust estimation and uncertainty quantification is important in other applications, as well, such as direction-of-arrival estimation [9, 10], channel estimation [11] and object tracking [12]. In these works, Bayesian approaches and covariance estimation are used to quantify the uncertainty or to build confidence regions. Non-Bayesian confidence region constructions were also applied for source localization in wireless sensor networks [13], while [14] applied least squares wavelet spectrogram (LSWS) to analyze global navigation satellite systems (GNSS) time series data and built confidence level surfaces to quantify the uncertainty.
@@
To overcome the issue of only asymptotically guaranteed confidence regions, a substantial amount of recent research in system identification focus on approaches with non-asymptotic guarantees [15]. One of the most common techniques is to derive Probably Approximately Correct (PAC) bounds for an estimate, such as LS, using the theory of concentration inequalities, which then induce confidence regions.
@@
Several works analyzed (closed-loop) dynamical systems, where the sample complexity of transition matrix estimations were studied for fully observable [16] and also non-observable state space settings [17, 18]. Even unstable linear dynamical systems were studied in [19]. A signal processing focused approach was investigated in [20], where a low-rank and structured sparse high-dimensional vector autoregressive (VAR) system was analyzed, and a new estimation technique was proposed with non-asymptotic probabilistic upper bounds on the estimation error. In the aforementioned works, there are strong assumptions on the noises, namely that they follow some specific distribution, in most cases they are Gaussian. Even the least restrictive results assume subgaussianity [21, 22], which covers a wide range of possible noises where the tail probabilities decay at least as fast as for the Gaussian distribution. Another limitation is that the confidence regions heavily rely on specific constants (hyper-parameters), which are coming from the assumptions on the system setting and the distributions of the inputs and the outputs. In practice, these hyper-parameters are usually unknown and only estimates can be used or bounds based on domain specific knowledge of experts. For finite impulse response (FIR) systems, PAC bounds for the LS estimate were given in [23], under a centered subgaussian noise assumption. Despite the relaxed noise assumption compared to the Gaussian case, building confidence regions using this results would still rely on the variance proxy of the subgaussian distribution.
@@
The fundamental construction of SPS is the indicator function, which is a hypothesis test that checks whether a given parameter is included in the confidence region. Using the indicator function to determine all the points of the confidence set would be computationally demanding, therefore in [31] an ellipsoidal outer approximation (EOA) algorithm was proposed, which gives a compact representation of the SPS region. An important feature of the constructed ellipsoid is that its center and its shape matrix coincide with that of the confidence ellipsoid constructed using the classical asymptotic theory, only its radius is different. This radius, which ensures finite sample guarantees, can be computed in polynomial time, by solving semidefinite optimization problems. The EOA construction of SPS was later generalized to ARX [35] and closed-loop state space models [34], as well.
Theoretical guarantees besides the exact coverage were rigorously proven for the acceptance region of the SPS indicator function, such as uniform strong consistency [36], which means that the SPS regions almost surely shrink around the true parameters, as the sample size increases. However, this result is asymptotic, and the finite sample performance of SPS remained an open question until recently. In our previous work [37], we have investigated the sample complexity of SPS for linear regression, i.e., we have derived non-asymptotic PAC bounds for the volumes of the acceptance regions of the SPS indicator function (hypothesis test). We have also shown that the sizes of the SPS indicator regions decrease at an optimal rate $\mathcal{O}(1/\sqrt{n})$, where $n$ is the sample size.
In this work we provide a finite sample analysis of the EOA of SPS for general linear regression problems with exogenous regressors, e.g., for FIR systems. The main contributions are:
  1. 1.
    We prove high probability upper bounds on the sizes of SPS EOA confidence ellipsoids. Our analysis builds on some results from the sample complexity analysis of standard SPS regions [37], however, we analyze the finite sample properties of a convex semidefinite program, which requires a significantly different approach than deriving concentration inequalities for the indicator function based construction of standard SPS.
  2. 2.
    We also show that the obtained bounds are “close” to the bounds of the original SPS region and that they also decrease at the optimal rate, that is, $\mathcal{O}(1/\sqrt{n})$.
  3. 3.
    Extensive simulation experiments are presented, as well. We demonstrate the difference between our theoretical bounds and the empirical sizes of the regions, furthermore, we show comparisons with the PAC bounds of [23] and set membership ellipsoids.
We emphasize that although our analysis use distribution and regressor dependent parameters, the SPS EOA does not require these parameters, and it can be applied under very general assumptions, unlike standard non-asymptotic solutions, e.g., based on PAC bounds or set membership ellipsoids.
[…]
@@ II-B Assumptions
An (integrable) random variable $W$ is $\sigma$-subgaussian[38, Definition 2.2], if for all $\lambda\in\mathbb{R},$ there is a $\sigma>0$, such that
@@
Note that the quantity $\sigma^{2}$ is often referred to as the variance proxy. Naturally, any Gaussian random variable with variance $\sigma^{2}$ is $\sigma$-subgaussian. Furthermore, any distribution with a bounded support on $[a,b]$ is subgaussian with $\sigma=(b-a)/2$ [38, Example 2.4]. Standard examples of such distributions include the uniform, the triangular and the beta distributions.
@@
A4 applies the definition of coherence from [39, Definition 1.2] along with the facts that $\text{range}(\Phi_{n})=\text{range}(\Phi_{{\scriptscriptstyle Q},n})$ and $\Phi_{{\scriptscriptstyle Q},n}\Phi_{{\scriptscriptstyle Q},n}^{\mathrm{T}}$ is an orthogonal projection onto $\text{range}(\Phi_{n})$. A consequence of this is that $1\leq\mu(\Phi_{n})\leq n/d$ [39]. This assumption provides an upper bound on how much of the excitation can be “concentrated” to a single time step, i.e., onto a particular regressor $\varphi_{t}$. In the presented theorems A4 will be assumed alongside with A2, hence, the thin QR-decomposition of $\Phi_{n}=\Phi_{{\scriptscriptstyle Q},n}\Phi_{{\scriptscriptstyle R},n}$ is unique, since matrix $\Phi_{{\scriptscriptstyle R},n}$ is full rank.
@@
The SPS EOA algorithm, presented as Algorithm 3, requires much milder assumptions on the noises and regressors. The introduced stronger assumptions are needed to analyze the size of the constructed confidence ellipsoids for any finite sample size and almost all realizations of the regressor vectors.
@@ III-A The SPS-Indicator algorithm
$f(\delta)\,\doteq\,\begin{cases}\sigma\hskip 0.85358pt(\hskip 0.85358pt8\hskip 1.42262ptd\ln^{\frac{1}{2}}(\tfrac{4}{\delta})+d)^{\frac{1}{2}}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq{1},\\[2.84526pt] \sigma\left(8\ln(\tfrac{4}{\delta})+d\right)^{\frac{1}{2}}&0<\delta<4e^{-(nd\lambda_{0})^{2}},\\ \end{cases}$
@@ IV Sample complexity of the
The SPS ellipsoid gives a compact representation of the confidence region, and compared to applying the SPS-Indicator, it can be used much more easily in various applications, from robust optimization to risk management. It is also significantly less demanding from a computational point of view, as it can be computed in polynomial time. The following theorem formalizes the sample complexity of these ellipsoids, i.e., it shows how their radii shrink as the sample size increases.
@@
By expanding $\left\lVert S_{0}(\theta)\right\rVert^{2}$ as in [31, Section VI.], we get that
@@
Then, by introducing $z\doteq\bar{R}_{n}^{\frac{1}{2}}(\theta-\hat{\theta}_{n})$, the optimization problem (16) can be reformulated as [31, Section VI.]
@@
Note that in [41, Theorem 1] it has been shown that the SPS confidence regions are bounded, if both the perturbed and the unperturbed regressors span the whole space. This means that assuming A2 and $m=2$, the SPS region is bounded if $A$ is positive definite, which is independent of $\{W_{t}\}$.
In case of a bounded SPS confidence region, the ellipsoid $\mathcal{E}_{1}=\{\theta:\|S_{0}(\theta)\|^{2}-\|S_{1}(\theta)\|^{2}\leq 0\}$ is compact and non-empty, since it is a closure of the bounded SPS-Indicator set $\{\theta:\|S_{0}(\theta)\|^{2}\prec_{\pi}\|S_{1}(\theta)\|^{2}\}$, and by construction of the SPS region $\|S_{0}(\hat{\theta}_{n})\|^{2}=0$, therefore the LSE is always included in $\mathcal{E}_{1}$. In Lemma 4 we show that if the SPS region is bounded, given $\{\varphi_{t}\}$ and $\{\alpha_{1,t}\}$, it almost surely holds that $\|S_{1}(\hat{\theta}_{n})\|^{2}\neq 0$. From the previous two properties, it follows that the optimization problem (16) is strictly feasible, i.e., Slater’s condition hold, if we work with bounded SPS confidence regions. Consequently, strong duality holds [42, Appendix B], therefore the value of the above optimization program (19) is equal to the value of its dual which can be formulated by using the Schur complement, as follows:
@@
The semidefinite optimization problem (27) is convex and it can be solved using standard convex optimization tools. Note that problem (27) was originally proposed by [31], but only for the case of bounded regions. Nevertheless, here we also allow and handle the case of unbounded regions, c.f. (28).
@@
$\gamma^{*}\doteq\begin{cases}\gamma_{0}^{*}&\text{if the SPS region is bounded,}\\ \infty&\text{if the SPS region is unbounded.}\\ \end{cases}$ (28)
@@
Lower bounds on variables $\gamma_{0}$ and $\xi$ (27) can be determined by the conditions for positive semidefiniteness of (generalized) Schur complements [42, Appendix A.5.5], that is
@@
To give a lower bound on $\xi$, we will reformulate the condition $\xi A^{\prime}_{0}(\xi)\succeq 0$. First, we rewrite $A$ as in [37]
@@
The proof of Lemma 3 is built upon the same ideas as the proof of [37, Lemma 4], but since our second term is different, we present it in Appendix A, for the sake of completeness.
@@
Let us fix an index $i\in[m-1],$ and introduce the “bad” event $\mathcal{B}_{i}$ that the $i$-th ellipsoid is larger than our bound, i.e.,
@@
From Lemma 3, we know that $\mathbb{P}(\mathcal{B}_{i})\leq\delta$, for all $i$. Then, by using a similar argument as in the proof of [40, Theorem 3], we can write the overall bad event, i.e., that there are at least $m-q$ ellipsoids for which our bound does not hold, as
@@
V Experimental Results
In this section we illustrate the sizes of the SPS ellipsoidal outer approximation (EOA) confidence regions via a series of numerical experiments. We consider the problem of robust parameter (coefficient) estimation of an FIR system, i.e., when a confidence ellipsoid is also provided for the given estimate. This problem is a fundamental task in many practical signal processing applications, such as channel estimation [43].
We compare the sizes of the SPS EOA regions with our theoretical bound given in Theorem 3, with the sizes of confidence ellipsoids based on the classical asymptotic [4] and novel finite sample PAC theory [23], furthermore, with set membership ellipsoids [25]. The asymptotic confidence regions are derived from the central limit theorem. For zero mean independent and identically distributed (i.i.d.) noise with a finite variance the LSE is asymptotically Gaussian for bounded regressors. As a consequence, the approximate (asymptotic) confidence regions can be constructed as
$\tilde{\mathcal{C}}_{n}\doteq\left\{\theta\in\mathbb{R}^{d}:(\theta-\hat{\theta}_{n})^{\mathrm{T}}\bar{R}_{n}(\theta-\hat{\theta}_{n})\leq\frac{\mu\hat{\sigma}_{n}^{2}}{n}\right\},$ (67)
where
$\hat{\sigma}^{2}_{n}\doteq\frac{1}{n-d}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\theta}_{n}),$ (68)
and $\varepsilon(\theta)$ is given in Algorithm 2. The probability that the true parameter $\theta^{*}$ is in the confidence region $\tilde{\mathcal{C}}_{n}$ is approximately $F_{\chi^{2}(d)}(\mu)$, where $F_{\chi^{2}(d)}$ is the cumulative distribution function of the $\chi^{2}$ distribution with $d$ degrees of freedom. Then, we have $\mathbb{P}(\theta^{*}\in\tilde{\mathcal{C}}_{n})\approx p$. We applied [23, Theorem 3.1] to obtain a finite sample PAC bound for the LSE error. In [23] a different (but equivalent) characterization of a subgaussian variable is used. We will refer to the following result as the DMR-bound.
Theorem 4.
Let $\varphi_{1},\dots,\varphi_{n}$ be the time shifted regressors of a linear FIR system with independent centered $\sigma_{\varphi}$-subgaussian entries of unit variance and assume that the noises $\{W_{t}\}_{t=1}^{n}$ are $\sigma_{w}$-subgaussian. Then, there is an absolute constant $C>0$ such that, for all $\nu\in(0,1)$, $\eta\in(0,2e^{-1}]$ and as long as
$n\geq C(\sigma_{\varphi}^{2}\vee\sigma_{\varphi}^{4})\frac{d\log(d\vee\eta^{-1})}{\nu^{2}},$ (69)
the LS estimate $\hat{\theta}_{n}$ satisfies with probability at least $1-\eta$:
$\|\hat{\theta}_{n}-\theta^{*}\|^{2}\leq\frac{2}{(1-\nu)^{2}}\frac{d}{n}(1+C\sigma_{\varphi}^{2}\sigma_{w}^{2}\log^{2}(2/\eta)).$ (70)
Our set membership ellipsoid implementation is based on the well-established algorithm of [44], which recursively builds the confidence ellipsoid around a weighted LSE and only updates it when the available data is informative, i.e., when a smaller global set membership ellipsoid can be built.
V-A Sample complexity for bounded noise
Figure 1: Size comparison of empirical $0.9$-level SPS ellipsoidal outer approximation, asymptotic and set membership confidence ellipsoids with the theoretical bound of Theorem 3 for a uniform noise distribution with $m=10$, $q=1$, $\delta=0.5$, $t_{0}=400$, $n=2000$ and $s=100$.
We consider a $2$-dimensional FIR system
$Y_{t}=b_{1}^{*}U_{t-1}+b_{2}^{*}U_{t-2}+W_{t},$ (71)
where $b_{1}^{*}=b_{2}^{*}=5$. In the first experiment we chose a bounded noise sequence to satisfy the assumption of set membership ellipsoids, hence, $\{W_{t}\}$ were i.i.d. uniform random variables with zero mean and boundaries $(-3,3)$. Consequently, the noises were symmetric, nonatomic and $\sigma$-subgaussian with $\sigma=\sqrt{3}$, which is the optimal variance proxy. The input process $\{U_{t}\}$ was autoregressive, given by
$U_{t}=aU_{t-1}+\sum_{i=1}^{5}c_{i}V_{t-i+1},$ (72)
with $a=0.7$, $c_{1}=1$, $c_{2}=0.775$, $c_{3}=0.55$, $c_{4}=0.325$, $c_{5}=0.1$, and $V_{t}$ is a Gaussian white random process with distribution $\mathcal{N}(0,1)$. From (71), the linear regression problem can be constructed as $\theta^{*}=[b_{1}^{*},b_{2}^{*}]^{\mathrm{T}}$ and $\varphi_{t}=[U_{t-1},U_{t-2}]^{\mathrm{T}}$.